### v1

@@ 1 The Agentic Leash: Growing Fuzzy Cognitive Map Dynamical Systems from Text
These causal FCM feedback dynamical systems form local fuzzy or partial causal rules from the sampled documents. This local causal structure in turn defines global equilibrium limit cycles that serve as scenario-like answers to causal what-if questions. They also define the very source of the FCM dynamical system’s agency – its evolving equilibrim limit cycles. This differs from ordinary feedforward agentic systems whose agency resides only in programmed commands. Mixing FCMs can give both richer learned causal knowledge bases and richer global equilibria. The extraction process is agentic [acharya2025agentic] or partially autonomous because the FCM’s evolving global equilibria command the LLM agents to fetch and process further text that then tends to change the commanding FCM equilibria. Related work uses an autoencoder-like mapping to convert FCMs to text and continue the reverberatory process [panda2025causal]. This bidirectional process keeps the FCM’s LLM agents on a type of flexible agentic leash.
Figure 1: A Large Language Model (LLM) extracts causal variables and their casual relationships out of a Wall Street Journal article from Henry Kissinger and colleagues about the promise of AI and then creates a Fuzzy Cognitive Map (FCM). The figure shows only 5 out of the 15 AI-extracted nodes and the directed weighted edges that connect them. The positive edges are in blue and the negative edges are in red. The figure highlights one of many feedback loops in the FCM in green. In this case: Growth of Human Cognition increases Human-AI Interactions but an increase in Human-AI Interactions decreases Human Cognition.
@@
Figure 4: Noun Extraction with an agentic LLM: LLM takes a paragraph of text as input and identifies the nouns, noun phrases, and pronouns present in the text. The figure on the right highlights the extracted nouns and noun phrases in red.
We need an LLM that can systematically extract causal variables and their causal relationships from the text without human supervision. These casual rules from the text should then predict the dynamics of the system described by the text even if the text does not mention the dynamics. FCMs can model this causal information as a weighted directed cyclic graph and also give qualitative equilibrium predictions from the equilibria of the approximating causal FCM [kosko1986differential, kosko1986fuzzy, kosko1988hidden, osoba2017fuzzy, ziv2018potential, glykas2010fuzzy, papageorgiou2013fuzzy, stach2010divide, kosko1988hidden, taber2007quantization].
@@
Figure 1 shows the simplest version of the FCM agentic set-up. A Large Language Model (LLM) takes the WSJ article “ChatGPT Heralds an Intellectual Revolution” by Henry Kissinger et alia as text input and extracts casual variables and causal relationships among those variables from the text. It then builds a Fuzzy Cognitive Map (FCM) out of those causal rules. Figure 1 shows only 5 of the 15 nodes from the extracted FCM. Figure 2 shows the complete FCM.
@@
Figure 14 mixes the extracted-FCMs from the same article with the same guided prompts but used 2 different LLM agents: Gemini-2.5 Pro and ChatGPT-4.1. The 24-node mixed FCM used equal mixing weights for both FCMs. Figure 13 mixes the corresponding edge matrices of the 2 mixture=-omponent FCMs using zero-padding and convex combination. Figure 15 shows the limit cycles of this final mixed FCM.
@@ 2 Extracting FCMs from Text and Related Work
3 Fuzzy Cognitive Map (FCM) Basics
Fuzzy Cognitive Maps (FCM) model causal dynamical systems as directed weighted graphs. The concept nodes describe the causal variables in the dynamical system and the directed edges describe the causal relationships between those nodes. FCMs allow feedback and therefore converge to non-trivial equilibria like limit cycles. FCMs model dynamical systems by approximating their underlying maps from inputs to equilibria.
@@ 3.1 Causal Edge Matrix
The directed edges of the FCM describe the causal relationships between concept nodes. An edge $e_{ij}$ from the $i^{\text{th}}$ concept node $C_{i}$ to the $j^{\text{th}}$ concept node $C_{j}$ says $C_{i}$ causes $C_{j}$”. The fuzzy edge-weights describe partial causality. The edge-weight $w_{ij}\in[-1,1]$ on the edge $e_{ij}$ gives the degree to which $C_{i}$ causes $C_{j}$:
@@
An $n\times n$ matrix $E$ describes all the directed weighted edges of a $n$-node FCM. The edge-weight $w_{ij}$ corresponds to the matrix element on the $i^{\text{th}}$ row and the $j^{\text{th}}$ column. The matrix element is zero if there is no edge between the corresponding node-pair.
@@ 3.2 Discrete Time Evolution of FCMs
A $n$-dimensional row vector $C(t)\in[0,1]^{n}$ describes the state of the FCM’s concept nodes at time $t$. The $i^{\text{th}}$ node is “active” at time $t$ if the $i^{\text{th}}$ component $C_{i}(t)$ of the state vector $C(t)$ is equal to or close to one. The $i^{\text{th}}$ node is “inactive” at time $t$ if $C_{i}(t)$ is equal to or close to zero. A node is partially active otherwise. The casual variables corresponding to the active nodes are present in the system and those corresponding to the inactive nodes are absent. The casual factors are partially present in the system if their corresponding nodes are partially active.
@@
Figure 9: A paragraph from from the WSJ article ChatGPT Heralds an Intellectual Revolution by Henry Kissinger et al.
@@ 3.3 FCM Equilibria
The set of all initial conditions $C(0)$ that lead to a given equilibrium describe the “basin of attraction” for that equilibrium. The FCM describes a map from these basins to their corresponding equilibrium attractors. The basins of the FCM’s equilibria partition the FCM’s input space. The FCM models a dynamical system by approximating its corresponding basin-to-equilibrium map.
@@ 4 LLM-Based FCM Extractor
Large Language Models (LLM) are good at parsing text and processing natural language. LLMs can extract causal information from text that describes dynamical systems and use it to model the underlying dynamical systems as FCMs. But LLMs hallucinate with unguided prompts. Their answer may also vary across multiple runs with the same prompt. The LLMs need guiding prompts that are designed to extract FCMs.
@@ 4.1 Step 1: Noun Extraction
Figure 13: FCMs’ edge matrices mix through zero-padding and convex combination described in equation (3). The edge matrix of the 15-node Gemini-extracted FCM is on the top-left and the edge matrix of the 20-node ChatGPT-extracted FCM is on the bottom-left. The edge-matrix of the 24-node equal-weight FCM-mixture is on the right. The bigger edge-weights are colored brighter in the matrices. The matrices on the left also highlight the zero-padded rows and columns of the mixture components.
@@ 5 Simulations
Figure 15: Limit cycle from the mixed Gemini-ChatGPT FCM. The time steps are along the $x$-axis and the nodes are along the $y$-axis. Active nodes are in yellow and the inactive nodes are in purple. The 2-step limit cycles oscillates between states “Human Knowledge” and “Human Understanding.”
@@
The same LLM then took a WSJ article “ChatGPT Heralds an Intellectual Revolution” by Henry Kissinger et al. as input. Figure 9 shows one paragraph from this article. An unguided-prompt to the LLM extracted a FCM with 12 nodes. Figure 10(a) shows the edge-matrix corresponding to this FCM. A guided prompt to the same LLM gave a 15-node FCM instead. Figure 2 shows this FCM and Figure 10(b) gives its edge matrix. This 15-node FCM converged to a 4-step limit cycle. Figure 12 shows this limit cycle.
@@
OpenAI’s ChatGPT-4.1 LLM agent repectively produced a 24-node FCM and a 20-node FCM based on the same unguided and guided prompts. Figure 11 shows the edge matrices correspoonding to these FCMs. The 15-node Gemini-FCM and the 20-node ChatGPT-FCM mixed with equal mixing weights to give the final 24-node mixed FCM. Figure 14 shows this FCM-mixture. Figure 13 shows its edge-matrix while Figure 15 shows its limit cycle.

### v2

@@ 1 The Agentic Leash: Growing Fuzzy Cognitive Map Dynamical Systems from Text
These causal FCM feedback dynamical systems form local fuzzy or partial causal rules from the sampled documents. This local causal structure in turn defines global equilibrium limit cycles that serve as scenario-like answers to causal what-if questions. They also define the very source of the FCM dynamical system’s agency – its evolving equilibrium limit cycles. This differs from ordinary feedforward agentic systems whose agency resides only in programmed commands. Mixing FCMs can give both richer learned causal knowledge bases and richer global equilibria. The extraction process is agentic [acharya2025agentic] or partially autonomous because the FCM’s evolving global equilibria command the LLM agents to fetch and process further text that then tends to change the commanding FCM equilibria. Related work uses an autoencoder-like mapping to convert FCMs to text and continue the reverberatory process [panda2025causal]. This bidirectional process keeps the FCM’s LLM agents on a type of flexible agentic leash.
Figure 1: A Large Language Model (LLM) extracts causal variables and their causal relationships out of a Wall Street Journal article from Henry Kissinger and colleagues about the promise of AI and then creates a Fuzzy Cognitive Map (FCM). The figure shows only 5 out of the 15 AI-extracted nodes and the directed weighted edges that connect them. The positive edges are in blue and the negative edges are in red. The figure highlights one of many feedback loops in the FCM in green. In this case: Growth of Human Cognition increases Human-AI Interactions but an increase in Human-AI Interactions decreases Human Cognition.
@@
Figure 4: Noun Extraction with an agentic LLM: The LLM takes a paragraph of text as input and identifies the nouns, noun phrases, and pronouns present in the text. The figure on the right highlights the extracted nouns and noun phrases in red.
We need an LLM that can systematically extract causal variables and their causal relationships from the text without human supervision. These causal rules from the text should then predict the dynamics of the system described by the text even if the text does not mention the dynamics. FCMs can model this causal information as a weighted directed cyclic graph and also give qualitative equilibrium predictions from the equilibria of the approximating causal FCM [kosko1986differential, kosko1986fuzzy, kosko1988hidden, osoba2017fuzzy, ziv2018potential, glykas2010fuzzy, papageorgiou2013fuzzy, stach2010divide, kosko1988hidden, taber2007quantization].
@@
Figure 1 shows the simplest version of the FCM agentic set-up. An LLM takes the WSJ article “ChatGPT Heralds an Intellectual Revolution” by Henry Kissinger et alia as text input and extracts causal variables and causal relationships among those variables from the text. It then builds an FCM out of those causal rules. Figure 1 shows only 5 of the 15 nodes from the extracted FCM. Figure 2 shows the complete FCM.
@@
Figure 14 mixes the extracted-FCMs from the same article with the same guided prompts but used 2 different LLM agents: Gemini-2.5 Pro and ChatGPT-4.1. The 24-node mixed FCM used equal mixing weights for both FCMs. Figure 13 mixes the corresponding edge matrices of the 2 mixture-component FCMs using zero-padding and convex combination. Figure 15 shows the limit cycles of this final mixed FCM.
@@ 2 Extracting FCMs from Text and Related Work
3 FCM Basics
FCMs model causal dynamical systems as directed weighted graphs. The concept nodes describe the causal variables in the dynamical system and the directed edges describe the causal relationships between those nodes. FCMs allow feedback and therefore converge to non-trivial equilibria like limit cycles. FCMs model dynamical systems by approximating their underlying maps from inputs to equilibria.
@@ 3.1 Causal Edge Matrix
The directed edges of the FCM describe the causal relationships between concept nodes. An edge $e_{ij}$ from the $i^{\text{th}}$ concept node $C_{i}$ to the $j^{\text{th}}$ concept node $C_{j}$ says $C_{i}$ causes $C_{j}$”. The fuzzy edge-weights describe partial causality. The edge weight $w_{ij}\in[-1,1]$ on the edge $e_{ij}$ gives the degree to which $C_{i}$ causes $C_{j}$:
@@
An $n\times n$ matrix $E$ describes all the directed weighted edges of a $n$-node FCM. The edge weight $w_{ij}$ corresponds to the matrix element on the $i^{\text{th}}$ row and the $j^{\text{th}}$ column. The matrix element is zero if there is no edge between the corresponding node-pair.
@@ 3.2 Discrete Time Evolution of FCMs
A $n$-dimensional row vector $C(t)\in[0,1]^{n}$ describes the state of the FCM’s concept nodes at time $t$. The $i^{\text{th}}$ node is “active” at time $t$ if the $i^{\text{th}}$ component $C_{i}(t)$ of the state vector $C(t)$ is equal to or close to one. The $i^{\text{th}}$ node is “inactive” at time $t$ if $C_{i}(t)$ is equal to or close to zero. A node is partially active otherwise. The causal variables corresponding to the active nodes are present in the system and those corresponding to the inactive nodes are absent. The causal factors are partially present in the system if their corresponding nodes are partially active.
@@
Figure 9: A paragraph from the WSJ article ChatGPT Heralds an Intellectual Revolution by Henry Kissinger et al.
@@ 3.3 FCM Equilibria
The set of all initial conditions $C(0)$ that lead to a given equilibrium describes the “basin of attraction” for that equilibrium. The FCM describes a map from these basins to their corresponding equilibrium attractors. The basins of the FCM’s equilibria partition the FCM’s input space. The FCM models a dynamical system by approximating its corresponding basin-to-equilibrium map.
@@ 4 LLM-Based FCM Extractor
LLMs are good at parsing text and processing natural language. They can extract causal information from text that describes dynamical systems and use it to model the underlying dynamical systems as FCMs. But LLMs hallucinate with unguided prompts. Their answer may also vary across multiple runs with the same prompt. The LLMs need guiding prompts that are designed to extract FCMs.
@@ 4.1 Step 1: Noun Extraction
Figure 13: FCMs’ edge matrices mix through zero-padding and convex combination described in equation (3). The edge matrix of the 15-node Gemini-extracted FCM is on the top-left and the edge matrix of the 20-node ChatGPT-extracted FCM is on the bottom-left. The edge matrix of the 24-node equal-weight FCM-mixture is on the right. The bigger edge weights are colored brighter in the matrices. The matrices on the left also highlight the zero-padded rows and columns of the mixture components.
@@ 5 Simulations
Figure 15: Limit cycle from the mixed Gemini-ChatGPT FCM. The time steps are along the $x$-axis and the nodes are along the $y$-axis. Active nodes are in yellow and the inactive nodes are in purple. The 2-step limit cycle oscillates between states “Human Knowledge” and “Human Understanding.”
@@
The same LLM then took a WSJ article “ChatGPT Heralds an Intellectual Revolution” by Henry Kissinger et al. as input. Figure 9 shows one paragraph from this article. An unguided-prompt to the LLM extracted a FCM with 12 nodes. Figure 10(a) shows the edge matrix corresponding to this FCM. A guided prompt to the same LLM gave a 15-node FCM instead. Figure 2 shows this FCM and Figure 10(b) gives its edge matrix. This 15-node FCM converged to a 4-step limit cycle. Figure 12 shows this limit cycle.
@@
OpenAI’s ChatGPT-4.1 LLM agent respectively produced a 24-node FCM and a 20-node FCM based on the same unguided and guided prompts. Figure 11 shows the edge matrices correspoonding to these FCMs. The 15-node Gemini-FCM and the 20-node ChatGPT-FCM mixed with equal mixing weights to give the final 24-node mixed FCM. Figure 14 shows this FCM-mixture. Figure 13 shows its edge matrix while Figure 15 shows its limit cycle.